The volume of a sample of gas is 650.0 ml at STP. What volume will the sample occupy at 40.0C and 950.0 torr?
1
Expert's answer
2012-09-27T11:47:36-0400
PV=nRT
V1=nRT1/P1=0.6500 l
V1=nR*273.15/100=0.650
l
nR=0.23796

V2=nRT2/P2 (950.0 torr is 126.67
kPa)
V2=nR*313.15/126.67
V2=0.27396*313.15/126.67=0.588289 l or 588.29 ml